Customers at Nunhold (Hatton Technology Park) will briefly lose internet connectivity later today when we replace a failing switch with new hardware. We will install the new switch in parallel with the existing hardware so that we can gracefully repatch customer connections. This will minimise downtime to around 15 seconds per customer, however as this… Read More »
During the time frame stated below we will be migrating the remaining customers at our Westwood cabinet from Planet hardware to ZyXEL. This work is being completed as the ZyXEL hardware is more reliable and offers additional features. Planned Work Ref: WN29012015 Planned Work Start (Local UK): 29/01/2015 21:00 Planned Work End (Local UK): 29/01/2015 22:00
We have been notified by the University of Warwick of some planned maintenance they’re undertaking. We do not anticipate a loss of service during the time frame stated below, however services fed from our University of Warwick POP should be considered ‘At-Risk’ during the maintenance window. Planned Work Ref: UOW30012015 Planned Work Start (Local UK): 30/01/2015… Read More »
WarwickNet engineers will be troubleshooting some reported low connection speeds on Pride Park Derby this afternoon. Some customers there may experience one or more brief losses of service between 15:00 – 17:00.
During the time frame stated below, we will be performing maintenance works at our IX London POP to improve network performance and allow for future expansion. Some sites will experience a loss of service, whilst this work takes place. Planned work Start: 28/01/15 – 23:00 Planned Work End: 29/01/15 – 02:00
20/1/2015 – 11pm – 1am During the above time slot we will be rolling out a refined VDSL2 profile to our DSLAMs. This will cause the line to resync causing approximately 30s of outage. All VDSL connected customers will be affected. Apologies for any inconvenience caused. Update: 20/1/2015 23:01 This maintenance is starting now. Update:… Read More »
We have been informed that the fire brigade are on-site at Lumonics House and are investigating reports of smoke in the roofspace. Currently our equipment remains powered, however this should be considered an “at risk” period, as it may be necessary for the power to be cut as a safety precaution. In that instance, all… Read More »
[UPDATE 16:00] Power has been restored to the building following the replacement of the meter. [UPDATE 15:00] Our generator is in place and is now powering the equipment on-site, which should provide adequate cover whilst the work is carried out to replace the faulty meter. [UPDATE: 14:20] EON have arrived unannounced on-site to replace the… Read More »